Smith Corp. bonds were issued 25 years ago. These bonds have a 15% required rate of return, mature in 11 years, and have a current market value of $995.10. Find the coupon rate. Round intermediate steps and your final answer to four decimals. Enter your answer in decimal format (EX: .XXXX).
Answer #1

Current Market Value of bonds is equal to the present value of all coupon payments and the principal amount

Let the coupon payment be x

995.10 = x*PVAF(15%, 11 years) + 1000*PVF(15%, 11 years)

995.10 = x*5.2337 +1,000*0.2149

x = 149.07235

Coupon Rate = 149.07235/1000 = 0.1491